OpenStudy (anonymous):

What is the value of g[f(5)] for the functions f(x) = 2x + 1 and g(x) = 2x - 5?

OpenStudy (anonymous):

2(2x+1)-5= 4x-3 = g(f(x)) putting x=5 will give you 17